OpenStudy (anonymous):

-2x + 4y = 48 x + 4y = 24

OpenStudy (anonymous):

we just answered a question like this for you, it is the same steps, just different numbers.

OpenStudy (anonymous):

subtract the second equation from the first -2x-x+4y-4y=48-24 -3x=24 x=24/-3 then substitute x into one of the equations to find y

OpenStudy (anonymous):



OpenStudy (anonymous):